OpenStudy (anonymous):

How to solve 2x+6=4x+8

OpenStudy (anonymous):

Divide whole equation by 2. x+3=2x+4 -x+3=4 -x=1

OpenStudy (anonymous):

you would subtract 4x from both sides leaving you with -2x+6=8 then you would subtract 6 from both sides giving you -2x=2 then you divde both sides by 2 the answer is -1

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

Algebra 1?

OpenStudy (anonymous):

Yeah I forgot how to solve those I needed a little refresh

OpenStudy (anonymous):

cool im in it to i just forgot what a multiplied by y/b was lol